OpenStudy (anonymous):

A boat has a speed of 5 mph in calm water. It takes the boat 3 hours to travel upstream but only 2 hours to travel the same distance downstream. Using the data in the table, what is the value of c, the speed of the current?

OpenStudy (***[isuru]***):

hi, when the boat travels upstream the current retards it motion and decrease the speed of boat but when boat travels downstream the current help the motion an increase the speed of boat .... this is the base to this problem lets assume the speed of current as "c" for the motion upstream... if the traveled distance is "S" distance = time x speed S = 3(5-c)----(1) for the motion downstream... as the distance is same S =2(5+c)-----(2) divide (1) by (2) s/s = 3(5-c)/2(5+c) 10 +2c = 15 - 3c 5c = 5 c = 1mph hope u get it!!